Circular No. 8330

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 T7 (LINEAR) Z. Sekanina, Jet Propulsion Laboratory, writes: “The sudden appear-

ance of a prominent antitail is the result of unusual coincidences. Due to projection effects the antitail could not develop before the end of March 2004, at which time the comet was too close to the sun to be observed from the ground (and not close enough to show up in the SOHO’s C3 coronagraphic images). From the orientation in the available images of Apr. 17–20, the dust contained in the antitail is found to have been ejected most probably between discovery (mid-October 2002) and ∼ 300 days before perihelion (the end of June 2003). Because the comet was not brightening dramatically between discovery (except perhaps in the first post-discovery days) and the 2003 conjunction with the sun, it is likely that the material in the antitail dates back to a fairly sudden onset of persistent dust production at or shortly before discovery (i.e., the comet was detected because it ‘turned on’ at about that time), or the comet’s dust-emission activity picked up rapidly in mid-2003, when it was behind the sun. The steeper light curve beginning in August 2003 can support this scenario. In either case, there was no antitail around 2003 Dec. 27 and none is expected to appear around 2004 June 26, the times of the earth’s transit across the comet’s orbital plane. In the intervening period of time, the position angle of the antitail and the angle it subtends with the prolonged radius vector (i.e., approximately with the plasma tail) are predicted to vary as follows (the first number in each interval refers to an onset time in late June 2003, the second to onset in mid-October 2002): Apr. 15, p.a. 42^{o}–45^{o}, angle with respect to the radius vector 142^{o}–145^{o}; 20, 32^{o}–37^{o}, 137^{o}–142^{o}; 25, 16^{o}–24^{o}, 125^{o}–133^{o}; May 5, 323^{o}–334^{o}, 80^{o}–91^{o}; 15, 291^{o}–295^{o}, 82^{o}–86^{o}; 20, 274^{o}– 278^{o}, 120^{o}–124^{o}; 25, 255^{o}–262^{o}, 130^{o}–137^{o}; 30, 236^{o}–251^{o}, 117^{o}–132^{o}; June 4, 197^{o}–232^{o}, 80^{o}–115^{o}; 9, 154^{o}–186^{o}, 37^{o}–69^{o}. The angular deviation from the radius vector dropping below 90^{o} implies the antitail’s disappearance. There are two maxima in this angle, one on Apr. 13, the second on May 21. Unless the dust production terminated early, more recent ejecta fill the entire sector between the antitail and the plasma tail. The visibility of this sector is affected by projection geometry and, to a degree, is anticorrelated with the antitail’s prominence. Also, in long-exposure images, which show the antitail to extend further out, its slight curvature may be noticed toward greater position angles (counterclockwise) farther from the comet, although this effect may be concealed by decreasing brightness.”

Visual total magnitudes and tail lengths by A. Pearce, Noble Falls, W. Australia (8×40 binoculars): Apr. 15.89 UT, 4.6, 2^{o}.5; 22.88, 4.2, 6^{o}.5.

2004 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green